# On real forms of Fermat hypersurfaces

### v1

@@ Abstract.
In this paper, we compute the number of real forms of Fermat hypersurfaces for degree $d\geq 3$ except the degree $4$ surface case, and give explicit descriptions of them.
@@ 1. Introduction
$\begin{cases}\left\lfloor\frac{n}{2}\right\rfloor+2&(d\equiv 1\pmod{2})\\ \frac{(n+3)(n+5)}{8}&(d\equiv 0,\ n\equiv 1\pmod{2})\\ \frac{(n+4)(n+6)}{8}&(d\equiv 0,\ n\equiv 0\pmod{2})\end{cases}$
We prove this theorem by computing $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$. In Section 3, we deal with the case $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, and in Section 4, we deal with the case $(n,d)=(1,3)$. In both cases, we have already known the automorphism group $\operatorname{\operatorname{Aut}}F^{n}_{d}$, as the former case is due to Shioda [Shi87], and the latter is due to the fact that $F^{1}_{3}$ is elliptic. So we will check the condition for an automorphism of $F^{n}_{d}$ to be a cocycle, compute a conjugation of a cocycle by an automorphism, and get this result.
We also give explicit descriptions of these real forms for the curve case in Theroem 5.1, and for the surface case in Theorem 5.2 in the way we can generalize it to higher dimensions, and give a topological description of a real locus of some real forms (in Theorem 5.3, and in particular, Theorem 5.4 for the curve case). We also show that $F_{4}^{2}$ has at least $4$ non-isomorphic real forms in Remark 5.5.
@@ 2. Preliminary about real forms
    Real structures $\rho$, $\rho^{\prime}$ on $X$ are equivalent if there exists a $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\rho^{\prime}=\varphi^{-1}\circ\rho\circ\varphi$.
@@
Let $X$ be a complex quasi-projective variety with a real structure $\rho$ on $X$. Then, there are bijections between
@@
  • cocycles $\alpha,\beta\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ are equivalent if and only if there exists $\varphi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\beta\circ\rho=\varphi^{-1}\circ(\alpha\circ\rho)\circ\varphi$, and we denote it by $\alpha\sim\beta$.
@@ 3. The number of real forms of Fermat hypersurfaces
Before thinking about real forms of Fermat hypersurfaces, we want to think about automorphisms of them. Clearly, there are two types of automorphisms when the base field is algebraically closed and of characteristic 00. An automorphism of the first type is the one with which we can identify $\sigma\in S_{n+2}$ defined as follows:
@@
An automorphism of the second type is the one denoted by $g(i,j)$ for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1 and defined as follows:
@@
with $\xi=\xi_{d}$ a primitive $d$-th root of unity.
Let $G^{n}_{d}$ be the subgroup of $\operatorname{\operatorname{Aut}}F^{n}_{d}$ generated by $g(i,j)$ for all $i,j\in\mathbb{Z}$. Then, the following result is known.
@@
$\operatorname{\operatorname{Aut}}F^{n}_{d,k}=G^{n}_{d}\cdot S_{n+2}$
where normal-$\cdot$ denotes a semidirect product.
In particular, $\operatorname{\operatorname{Aut}}F^{n}_{d,\mathbb{C}}=G^{n}_{d}\cdot S_{n+2}$.
Though this theorem is proved only in the curve case in that paper, the same argument works in higher dimensions as we can show that every automorphism is linear.
@@
Let $\tilde{\rho}\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1}$ be the complex conjugation, i.e., the $\mathbb{R}$-morphism defined by $\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X_{i}$ to $X_{i}$ for all $i$ and $c\in\mathbb{C}$ to $\bar{c}$. This is a real structure of $\mathbb{P}^{n+1}$, and it induces the real structure $\rho\colon F^{n}_{d}\to F^{n}_{d}$ on $F^{n}_{d}$.
@@
By straightforward calculations, for $i,j\in\mathbb{Z}$, $10\leq i\leq n+$ 1, $\sigma\in S_{n+2}$,
  • $g(i,j)\circ\sigma=\sigma\circ g(\sigma^{-1}(i),j)$.
  • $\sigma\circ\rho=\rho\circ\sigma$.
  • $g(i,j)\circ\rho=\rho\circ g(i,d-j)=\rho\circ g(i,j)^{-1}$.
The following result is the main step to prove the main theorem.
@@
    $\alpha=\mu\sigma\in\operatorname{\operatorname{Aut}}F^{n}_{d}$, with $\mu=g(0,p_{0})\cdots g(n+1,p_{n+1})\in G_{d}^{n}$, $\sigma\in S_{n+2}$, is a cocycle if and only if the following two conditions hold:
@@
    • $p_{i}\equiv p_{\sigma(i)}\pmod{d}$ if $i\in\operatorname{supp}\sigma$.
    Here, $\operatorname{supp}\sigma$ is the support of a permutation $\sigma$, i.e., $\operatorname{supp}\sigma:=\left\{i|\sigma(i)\neq i\right\}$.
@@
    Cocycles $\beta=\lambda\tau$, $\gamma$ with $\lambda\in G^{n}_{d}$ and $\tau\in S_{n+2}$ are equivalent if and only if there exist $\psi\in S_{n+2}$ and $q_{j}\in\mathbb{Z}$ for $10\leq j\leq n+$ 1 with $q_{j}\equiv q_{\sigma(j)}\pmod{2}$ for all $j$ and $q_{l}\equiv 0\pmod{2}$ for all $l\notin\operatorname{supp}\tau$ such that
    $\gamma=\psi^{-1}\cdot\left(g(0,q_{0})\cdots g(n+1,q_{n+1})\beta\right)\cdot\psi.$
@@
$(\mu\sigma\rho)^{2}$ $=\mu\sigma\mu^{-1}\sigma$
$=g(0,{p^{\prime}}_{0})\cdots g(n+1,{p^{\prime}}_{n+1})\cdot\sigma^{2}$
where
${p^{\prime}}_{i}=\begin{cases}0&(i\notin\operatorname{supp}\sigma)\\ p_{i}-p_{\sigma(i)}&(i\in\operatorname{supp}\sigma)\end{cases}$
Therefore, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and ${p^{\prime}}_{i}\equiv 0\pmod{d}$, and the result follows.
(2) Suppose $\beta=\lambda\tau$ be a cocycle and $\varphi=\nu\psi\in\operatorname{\operatorname{Aut}}F^{n}_{d}$ with $\lambda,\nu=g(0,r_{0})\cdots g(n+1,r_{n+1})\in G^{n}_{d},\ \tau,\psi\in S_{n+2}$. Then,
$\varphi^{-1}\circ\beta\rho\circ\varphi$
$=$ $(\psi^{-1}\nu^{-1})\cdot(\lambda\tau\rho)\cdot(\nu\psi)$
$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\tau\nu^{-1})\cdot\psi\rho$
$=$ $\psi^{-1}\cdot(\nu^{-1}\lambda\cdot g(\tau(0),d-r_{0})\cdots g(\tau(n+1),d-r_{n+1}))\cdot\tau\psi\rho$
$=$ $\psi\cdot(g(0,s_{0})\cdots g(n+1,s_{n+1})\cdot\lambda\tau)\cdot\psi\rho$
where
$s_{i}=\begin{cases}-2r_{i}&(i\notin\operatorname{supp}\tau)\\ -r_{i}-r_{\tau(i)}&(i\in\operatorname{supp}\tau)\end{cases}$
As $r_{i}$ varies for all integers, the result follows. ∎
Note that, as cocycles,
@@
for $m\in\mathbb{Z}$ by definition, and
$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{2i}&\\ &I_{r}\end{pmatrix},\ \begin{pmatrix}I_{r}&&\\ &0&1\\ &1&0\end{pmatrix}\sim\begin{pmatrix}I_{r}&&\\ &0&\xi^{j}\\ &\xi^{j}&0\end{pmatrix}$
for $i,j\in\mathbb{Z}$ by the proposition.
By this proposition, we can describe $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
@@
Suppose $d$ is odd. Then, any cocycles are equivalent to exactly one of the following forms:
[…]
@@
$\begin{pmatrix}1&\\ &I_{r}\end{pmatrix}\sim\begin{pmatrix}\xi^{i}&\\ &I_{r}\end{pmatrix}$
for all $i$, and the former statement follows from Proposition 3.3. Also, as any two of $H^{r}_{n}$’s are not equivalent,
@@
where we identify cocycles with their equivalent classes, and the latter statement follows. ∎
@@
Suppose $d$ is even. Then, any cocycles are equivalent to exactly one of the forms:
@@
Also, $F^{n}_{d}$ has exactly $\frac{(n+3)(n+5)}{8}$ real forms for $n$ odd, and $\frac{(n+4)(n+6)}{8}$ real forms for $n$ even.
In particular, if $d\geq 3$ is even, then $F^{1}_{d}$ has exactly $3$ real forms and $F^{2}_{d}$ ($d\not=4$) has $6$ real forms.
@@
The former statement is clear from proposition 3.3.
By the proposition, for $s,s^{\prime},t,t^{\prime}\in\mathbb{Z}$ with $s+t=s^{\prime}+t^{\prime}$, $K^{s,t}_{n}\sim K^{s^{\prime},t^{\prime}}$ if and only if there exists a permutation matrix $\sigma$ and $a_{1}$, $\ldots$, $a_{n+2}$ with $a_{1}\equiv\cdots\equiv a_{m}\pmod{2}$ such that
$K^{s,t}_{n}\cdot\left(\sigma^{-1}K^{s^{\prime},t^{\prime}}_{n}\sigma\right)^{-1}=\begin{pmatrix}\xi^{a_{1}}&&\\ &\ddots&\\ &&\xi^{a_{n+2}}\end{pmatrix}.$
This condition is equivalent to $s=s^{\prime}$ or $s=t^{\prime}$.
Therefore,
[…]
where we identify cocycles with their equivalent classes. From this, when $n=2k-1$ for $k\in\mathbb{Z}$, the number of real forms is

### v2

@@ Abstract.
In this paper, we compute the number of real forms of Fermat hypersurfaces for degree $d$ except the degree $4$ surface case, and give explicit descriptions of them.
@@ 1. Introduction
$\begin{dcases}\left\lfloor\frac{n}{2}\right\rfloor+2&(d\equiv 1\pmod{2})\\ \frac{(n+3)(n+5)}{8}&(d\equiv 0,\ n\equiv 1\pmod{2})\\ \frac{(n+4)(n+6)}{8}+1&(d\equiv 0,\ n\equiv 0\pmod{2})\end{dcases}$.
Also, the number of real forms of $F^{n}_{2,\mathbb{C}}$ is
$\begin{dcases}\frac{n+1}{2}+1&(n\equiv 1\pmod{2})\\ \frac{n}{2}+3&(n\equiv 0\pmod{2})\end{dcases}$.
We prove this theorem by computing $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$. In Section 3, we deal with the case $(n,d)\neq(1,3),(2,4)$, $d\geq 3$, in Section 4, we deal with the case $(n,d)=(1,3)$, and in Section 5, we deal with the case $d=2$. In every case, we have already known the automorphism group $\operatorname{\operatorname{Aut}}F^{n}_{d}$, as the first case is due to Shioda [Shi87], the second is due to the fact that $F^{1}_{3}$ is elliptic, and in the third case, $\operatorname{\operatorname{Aut}}F^{n}_{2}\cong\operatorname{PO}(n+2)$. So we will check the condition for an automorphism of $F^{n}_{d}$ to be a cocycle, compute a conjugation of a cocycle by an automorphism, and get this result.
We also give explicit descriptions of these real forms for the curve case in Theorem 6.2, and for the surface case in Theorem 6.3 in the way we can generalize it to higher dimensions, and give a topological description of a real locus of some real forms (in Theorem 6.4, and in particular, Theorem 6.5 for the curve case). We also show that $F_{4}^{2}$ has at least $4$ non-isomorphic real forms in Remark 6.6.
This paper is a refined version of my master thesis [S21] (in Japanese, unpublished) under Professor Keiji Oguiso, together with a correction of some mistakes in the preliminary version, which was kindly pointed out by Professor János Kollár.
@@ 2. Preliminary about real forms
    Real structures $\rho$, $\rho^{\prime}$ on $X$ are equivalent if there exists a $\phi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\rho^{\prime}=\phi^{-1}\circ\rho\circ\phi$.
@@
Let $X$ be a complex quasi-projective variety with a real structure $\rho$. Then, there are bijections between
@@
  • cocycles $\alpha,\beta\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ are equivalent if and only if there exists $\phi\in\operatorname{\operatorname{Aut}}_{\mathbb{C}}X$ such that $\beta\circ\rho=\phi^{-1}\circ(\alpha\circ\rho)\circ\phi$, and we denote it by $\alpha\sim\beta$.
@@ 3. The number of real forms of Fermat hypersurfaces
In order to study real forms of Fermat hypersurfaces, we first recall automorphisms of them. Clearly, there are two types of automorphisms when the base field is algebraically closed and of characteristic 0. An automorphism of the first type is the one with which we can identify $\sigma\in\mathfrak{S}_{n+2}$ defined as follows:
@@
An automorphism of the second type is the one denoted by $g(i,j)$ for $i,j\in\mathbb{Z}$, $0\leq i\leq n+1$ and defined as follows:
@@
with $\xi=\xi_{d}$ a primitive $d$-th root of unity. Also, we denote the product $g(0,j_{0})\cdot\cdots\cdot g(n,j_{n})$ by $D\left(j_{0},\ldots,j_{n}\right)$.
Let $G^{n}_{d}$ be the subgroup of $\operatorname{\operatorname{Aut}}F^{n}_{d}$ generated by $g(i,j)$ for all $i,j\in\mathbb{Z}$. Then, the following result is known:
@@
$\operatorname{\operatorname{Aut}}F^{n}_{d,k}=G^{n}_{d}\cdot\mathfrak{S}_{n+2}$
where $\cdot$ denotes a semidirect product.
In particular, $\operatorname{\operatorname{Aut}}F^{n}_{d,\mathbb{C}}=G^{n}_{d}\cdot\mathfrak{S}_{n+2}$.
Though this theorem is proved only in the curve case in that paper, the same argument works in higher dimensions, as we can show that every automorphism is linear.
@@
Let $\tilde{\rho}\colon\mathbb{P}^{n+1}\to\mathbb{P}^{n+1}$ be the complex conjugation, i.e., the $\mathbb{R}$-morphism defined by $\mathbb{C}[X]\to\mathbb{C}[X]$ that sends $X_{i}$ to $X_{i}$ for all $i$ and $c\in\mathbb{C}$ to $\overline{c}$. This is a real structure of $\mathbb{P}^{n+1}$, and it induces the real structure $\rho\colon F^{n}_{d}\to F^{n}_{d}$ on $F^{n}_{d}$.
@@
The following relations hold for $D\left(j_{0},\ldots,j_{n+1}\right)$, $\sigma\in\mathfrak{S}_{n+2}$, and $\rho$.
$\sigma\cdot D\left(j_{0},\ldots,j_{n+1}\right)\cdot\sigma^{-1}=D\left(j_{\sigma(0)},\ldots,j_{\sigma(n+1)}\right)$
$\sigma\cdot\rho=\rho\cdot\sigma$
$D\left(j_{0},\ldots,j_{n+1}\right)\cdot\rho=\rho\cdot D\left(-j_{0},\ldots,-j_{n+1}\right)=\rho\cdot D\left(j_{0},\ldots,j_{n+1}\right)^{-1}.$
The following result is the main step in proving the main theorem.
@@
    $\alpha=\mu\sigma\in\operatorname{\operatorname{Aut}}F^{n}_{d}$, with $\mu=D\left(p_{0},\ldots,p_{n+1}\right)\in G_{d}^{n}$, $\sigma\in\mathfrak{S}_{n+2}$, is a cocycle if and only if the following conditions are satisfied:
@@
    • $p_{i}\equiv p_{\sigma(i)}\pmod{d}$ for any $0\leq i\leq n+1$ if $\sigma$ has a fixed point, i.e., $\sigma(k)=k$ for some $k$.
    • $p_{i}-p_{\sigma(i)}\equiv p_{j}-p_{\sigma(j)}\equiv 0,\dfrac{d}{2}\pmod{d}$ for any $0\leq i,j\leq n+1$ if $\sigma$ has no fixed point, i.e., $\sigma$ is of type $2^{\frac{n+2}{2}}$.
@@
    Cocycles $\beta=\lambda\tau$ with $\lambda\in G^{n}_{d}$ and $\tau\in\mathfrak{S}_{n+2}$ and $\gamma$ are equivalent if and only if there exist $\psi\in\mathfrak{S}_{n+2}$ and $r_{j}\in\mathbb{Z}$ for $0\leq j\leq n+1$ with $r_{j}=r_{\sigma(j)}$ for all $j$ and $r_{l}\equiv 0\pmod{2}$ for all $l\notin\operatorname{supp}\tau$ such that
    $\gamma=\psi^{-1}\cdot\left(D\left(r_{0},\ldots,r_{n+1}\right)\cdot\beta\right)\cdot\psi.$
    Here, $\operatorname{supp}\tau$ is the support of a permutation $\tau$, i.e., $\operatorname{supp}\tau:=\left\{i|\tau(i)\neq i\right\}$.
@@
$(\mu\sigma\rho)^{2}$ $=\mu\sigma\rho\mu\sigma\rho=\mu\sigma\mu^{-1}\sigma$
$=D\left(p_{0},\ldots,p_{n+1}\right)\cdot D\left(-p_{\sigma(0)},\ldots,-p_{\sigma(n+1)}\right)\cdot\sigma^{2}$
$=D\left(p_{0}-p_{\sigma(0)},\ldots,p_{n+1}-p_{\sigma(n+1)}\right)\cdot\sigma^{2}$
Therefore, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $(\mu\sigma\rho)^{2}=D\left(N,\ldots,N\right)$ for some $N\in\mathbb{Z}$.
When $\sigma(i)=i$, as $p_{i}-p_{\sigma(i)}=0$, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and $p_{j}-p_{\sigma(j)}\equiv 0\pmod{d}$ for any $j$.
When $\sigma$ has no fixed point, $(\mu\sigma\rho)^{2}=\operatorname{id}$ is equivalent to $\sigma^{2}=\operatorname{id}$ and $p_{0}-p_{\sigma(0)}\equiv\cdots\equiv p_{n+1}-p_{\sigma(n+1)}\pmod{d}$. In this case, for $0\leq i\leq n+1$,
$p_{i}-p_{\sigma(i)}\equiv p_{\sigma(i)}-p_{\sigma^{2}(i)}\equiv p_{\sigma(i)}-p_{i}\pmod{d}$
as $\sigma^{2}=\operatorname{id}$. Therefore, in this case, $p_{i}-p_{\sigma(i)}\equiv 0,\frac{d}{2}$ as $\sigma(i)\neq i$, and this is exactly what we want to show.
(2) Suppose $\beta=\lambda\tau$ be a cocycle and $\phi=\nu\psi\in\operatorname{\operatorname{Aut}}F^{n}_{d}$ with $\lambda,\nu=D\left(q_{0},\ldots,q_{n+1}\right)\in G^{n}_{d},\ \tau,\psi\in\mathfrak{S}_{n+2}$. Then,
$\phi^{-1}\cdot\beta\rho\cdot\phi$
$=$ $\left(\psi^{-1}\nu^{-1}\right)\cdot\left(\lambda\tau\rho\right)\cdot\left(\nu\psi\right)$
$=$ $\psi^{-1}\cdot\left(\nu^{-1}\lambda\tau\nu^{-1}\right)\cdot\psi\rho$
$=$ $\psi^{-1}\cdot\left(\nu^{-1}\lambda\cdot D\left(-q_{\tau(0)},\ldots,-q_{\tau(n+1)}\right)\tau\right)\cdot\psi\rho$
$=$ $\psi^{-1}\cdot\left(D\left(q_{0}+q_{\tau(0)},\ldots,q_{n+1}+q_{\tau(n+1)}\right)^{-1}\lambda\tau\right)\cdot\psi\rho$
As $q_{i}$’s vary for all integers, the result follows. ∎
Remark 3.4.
By Proposition 3.3, any cocycles in $\operatorname{\operatorname{Aut}}F^{n}_{d}$ are equivalent to one of the following forms:
$\begin{pmatrix}\xi^{a_{0}}&&\\ &\ddots&\\ &&\xi^{a_{n+1}}\end{pmatrix}\cdot\begin{pmatrix}I_{r}&&&&&\\ &0&1&&&\\ &1&0&&&\\ &&&\ddots&&\\ &&&&0&1\\ &&&&1&0\end{pmatrix}$
with $I_{r}$ the identity matrix of size $r$, $a_{i}\in\mathbb{Z}$, and $0\leq r\leq n+2$. Throughtout this paper, blank entries in matrices mean 0. Also, as cocycles,
@@
for $m\in\mathbb{Z}$ by definition, and in $\operatorname{\operatorname{Aut}}\left(F_{d}^{n}\right)$
$\begin{pmatrix}1&\\ &A\end{pmatrix}\sim\begin{pmatrix}\xi^{2i}&\\ &A\end{pmatrix},\ \begin{pmatrix}B&&\\ &0&1\\ &1&0\end{pmatrix}\sim\begin{pmatrix}B&&\\ &0&\xi^{j}\\ &\xi^{j}&0\end{pmatrix}$
for $i,j\in\mathbb{Z}$ and any matrices $A,B$ by Proposition 3.3.
Now, we can describe $H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})$.
@@
Suppose $d$ is odd. Then, any cocycles in $\operatorname{\operatorname{Aut}}F^{n}_{d}$ are equivalent to exactly one of the following forms:
$H^{r}_{n}:=\begin{pmatrix}I_{r}&&&&&\\ &0&1&&&\\ &1&0&&&\\ &&&\ddots&&\\ &&&&0&1\\ &&&&1&0\end{pmatrix}\ (r\in\mathbb{Z},\ 0\leq r\leq n+2).$
@@
$\begin{pmatrix}1&\\ &A\end{pmatrix}\sim\begin{pmatrix}\xi^{i}&\\ &A\end{pmatrix}$
for all $i$ and matrices AA. Therefore, when $d$ is odd, cocycles $\mu\sigma$, $\lambda\tau$ with $\mu,\lambda\in G_{d}^{n}$ and $\sigma,\tau\in\mathfrak{S}_{n+2}$ are equivalent if and only if the types of $\sigma$ and $\tau$ agree by Proposition 3.3 and Remark 3.4. Therefore,
@@
where we identify cocycles with their equivalent classes, and the statement follows. ∎
@@
Suppose $d$ is even. Then, any cocycles in $\operatorname{\operatorname{Aut}}F^{n}_{d}$ are equivalent to exactly one of the forms:
@@
$L_{n}:=\begin{pmatrix}0&1&&&\\ -1&0&&&\\ &&\ddots&&\\ &&&0&1\\ &&&-1&0\end{pmatrix}\ (n\equiv 0\pmod{2})$
Also, $F^{n}_{d}$ has exactly $\frac{(n+3)(n+5)}{8}$ real forms for $n$ odd, and $\frac{(n+4)(n+6)}{8}+1$ real forms for $n$ even.
In particular, if $d\geq 3$ is even, then $F^{1}_{d}$ has exactly $3$ real forms and $F^{2}_{d}$ ($d\not=4$) has $7$ real forms.
@@
By Proposition 3.3 and Remark 3.4, any cocycles are equivalent to $L_{n}$ or
$d_{r}(a_{1},\ldots,a_{r}):=\begin{pmatrix}\xi^{a_{1}}&&&&&&&\\ &\ddots&&&&&&\\ &&\xi^{a_{r}}&&&&&\\ &&&0&1&&&\\ &&&1&0&&&\\ &&&&&\ddots&&\\ &&&&&&0&1\\ &&&&&&1&0\end{pmatrix}$
with $a_{i}=0,1$ for each $i$. Then, by Proposition 3.3, $\alpha:=d_{r}(a_{1},\ldots,a_{r})$ is equivalent to $\beta:=d_{r}(b_{1},\ldots,b_{r})$ if and only if $\beta\sigma^{-1}\alpha^{-1}\sigma=d_{r}(c_{1},\ldots,c_{r})$ with $\sigma\in\mathfrak{S}_{n+2}$ and $c_{i}\equiv c_{j}\pmod{2}$ for all $i,j$, which is equivalent to say that $n(\alpha)=n(\beta)$ or $n(\alpha)+n(\beta)=r$ where $n(\alpha)=\#\left\{i|a_{i}=1\right\}$. Also, $d_{r}(a_{1},\ldots,a_{r})$ and $d_{s}(b_{1},\ldots,b_{s})$ are not equivalent if $r\neq s$ by Propositon 3.3. Therefore,
$H^{1}(\operatorname{Gal}(\mathbb{C}/\mathbb{R}),\operatorname{\operatorname{Aut}}F^{n}_{d})\setminus\{L_{n}\}=\left\{K^{s,t}_{n}|s+t\leq n+2,\ 0\leq s\leq t\right\}$
where we identify cocycles with their equivalent classes. From this, when n=2k1n=2k-1 for $k\in\mathbb{Z}$, the number of real forms is